Calculate the population of Kumasi for 2021 if the growth rate is 3% per annum and the death rate is 1%. The impact of the opening of the borders to other ecowas countries brings in 4% immigrants from neighbouring other countries. Residents leaving the Kumasi for greener pastures is also 0.05%


1
Expert's answer
2021-02-24T07:03:47-0500
"3-1+4-0.05=5.95"

The metro area population of Kumasi in 2020 was 3,348,000.


"3348000\times\dfrac{100+5.95}{100}=3547206"

The metro area population of Kumasi in 2020 will be 3,547,206.